If a system has an infinite number of solutions, use set-builder notation to write the solution set. If a system has no solution, state this.
Solve using the substitution method.
3y + x = 4,
x = 2y - 1 Answer by hmmimstuck(1) 
 Plug the second equation into the first equation:
3y+(2y-1)=4
5y-1=4
5y=5
y=1
Now plug your answer into the second equarion:
x=2(1)-1
x=2-1
X=1
To check your work plug both answers into the same equation:
3(1)+(1)=4
3+1=4
4=4